OpenStudy (anonymous):

Please help...I don't know what to do or even where to start. Given the equation (x+3)^2+y^2=16 (a) find the center and the radius (b) graph (c) find the intercepts, if any.

OpenStudy (anonymous):

foil first

OpenStudy (anonymous):

(x+3)(x+3)

OpenStudy (anonymous):

x=-3 and y=0 center radius =4

OpenStudy (asnaseer):

the general equation of a circle is:\[(x-a)^2+(y-b)^2=r^2\]where (a,b) is the coordinate of the center of the circle, and r is its radius.

OpenStudy (anonymous):



OpenStudy (anonymous):

x when x=0 y= +-4

OpenStudy (asnaseer):

so if you rewrite your equation:\[(x+3)^2+y^2=16\]in the form you get:\[(x-(-3))^2+(y-0)^2=4^2\]from which you can see the center must be at (-3, 0) and the radius must be 4 as @cinar said above.

OpenStudy (anonymous):

I am sorry I made a mistake when x=0 y=+-sqrt7

OpenStudy (anonymous):